OpenStudy (anonymous):

write the exponetial function y=ab^x for a graph that includes (2,24) and (3,48)

OpenStudy (anonymous):

\[ab^2=24\] \[ab^3=48\] so \[\frac{ab^3}{ab^2}=b=\frac{48}{24}=2\]

OpenStudy (anonymous):

\[a\times 2^2=24\] \[4a=24\] \[a=6\] and we can check that \[6\times 2^3=6\times 8 =48\] so it works.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh so the answer is \[y=6\times 2^x\]

OpenStudy (anonymous):

but whats first the 2 then 6?

OpenStudy (anonymous):

first we solved for b, so b = 2. then we solved for a and a = 6 you understand that only the 2 is being raised to the power of x. \[y = 6\times 2^x\]